OpenStudy (anonymous):

HELP PLEASE Solve the system using the addition method. 5x - 2y = 3 2x - y = 0 What is the y-value in the solution?

OpenStudy (anonymous):

try multiplying the 2nd equation by -2 then adding the 2 equations together

OpenStudy (anonymous):

ok lets see

OpenStudy (anonymous):

-2?

OpenStudy (anonymous):

-2(2x - y) = -2(0) what do you get?

OpenStudy (anonymous):

-4x + 2y = 0

OpenStudy (anonymous):

ok now add that to the first equation

OpenStudy (anonymous):

1x = 0

OpenStudy (anonymous):

er, 3+0 =?

OpenStudy (anonymous):

3

OpenStudy (anonymous):

so if you add up the converted 2nd equation to the first, you should get 1x =3 because the 2y cancel out

OpenStudy (anonymous):

yeah i see 2 cancels ....just didnt get that .......ohh i see what i did okok

OpenStudy (anonymous):

ok, since you have x can you solve for y?

OpenStudy (anonymous):

should i plug it in 5x -2y = 3

OpenStudy (anonymous):

its 6 (:

OpenStudy (anonymous):

you could plug it into the other equation you should get the same answer

OpenStudy (anonymous):

it a good way of checking, cause if you didnt get hte same answer, then one of your numbers is wrong

OpenStudy (anonymous):

i checked